# Capital Asset Pricing Model with Size Factor and Normalizing by Volatility Index

### v1

@@ Abstract.
The Capital Asset Pricing Model (CAPM) relates a well-diversified stock portfolio to a benchmark portfolio. We insert size effect in CAPM, capturing the observation that small stocks have higher risk and return than large stocks, on average. Dividing stock index returns by the Volatility Index makes them independent and normal. In this article, we combine these ideas to create a new discrete-time model, which includes volatility, relative size, and CAPM. We fit this model using real-world data, prove the long-term stability, and connect this research to Stochastic Portfolio Theory.
@@ 1. Introduction
Idea 2: Size Effect. Small stocks, on average, have higher volatility but also higher returns than large stocks. This feature implies long-term stability: Well-diversified stock portfolios stay together and not split into several clouds. We combined these two ideas in the previous article by the second author [11]. This article is a continuation of that work. Such long-term stability is of interest in Stochastic Portfolio Theory, which constructs portfolios as functions of market weights, utilizing the observation that small stocks have higher risk and return.
@@
In this article, we continue research of [11] by inserting Idea 3 in this setting, thus combining all three ideas. We create a new discrete-time model, and fit it using real market data from Kenneth French’s data library and Federal Reserve Economic Data web site. We then state and prove long-term stability of this market model, and interpret this for Stochastic Portfolio Theory. Unlike [11], we consider only discrete time models in this article.
In Section 2, we provide a comprehensive motivation of proposed models and historical review. Section 3 is devoted to data description and statistical fit of these models. In Section 4, we state and prove long-term stability (ergodicity) result for this model: Theorem 1, and analyze sufficient conditions for Theorem 1 to hold. Finally, in Section 5, we interpret this in terms of Stochastic Portfolio Theory. The Appendix contains a short discussion of stability conditions for the case where volatility is constant. We have already discussed this in our previous article [11] but only for continuous time. Thus it left the gap in our research. Here, we fill it.
@@ 2.2. Size and value
Including a few factors (for example size and value) would enrich (3). In particular, size factor is related to $\beta$ as follows: Well-diversified portfolios of small stocks have equity premia $P$ closely correlated with that $P_{0}$ of large stock benchmark S&P 500. This simple linear regression of $P$ vs $P_{0}$ has very large $R^{2}$ but $\beta$ which is slightly larger as $1$. This can be done, for example, with exchange-traded funds tracking small, mid, and large (=benchmark) stock indices, see [11, Appendix]. The $\beta$ for mid-cap index is 1.15, and for small-cap index is 1.27. A natural idea then is to model $\beta$ as a function of a portfolio size relative to benchmark. We can try the same for $\alpha$, although in [11, Appendix] we have $\alpha$ is not significantly different from zero for both small-cap and mid-cap indices. See more on the size effect in [18], and the discussion in [1, Chapter 8].
@@ 2.3. CAPM-based model with size as factor
In fact, in [11] we have more general conditions: $\alpha$ and $\beta$ are general functions of $C$. Here, we focus only on linear functions, see [11, Example 1]. Analysis for small-cap and mid-cap funds above shows that $a\approx 0$ but $b<0$. Rewrite (3) as follows by plugging there (5):
@@ 2.4. Stochastic volatility
However, there is a drawback in our modeling from [11]. We fit regressions (6) and (7) using real-world monthly data from 1926 taken from Kenneth French’s Dartmouth College Financial Data Library. Residuals $\varepsilon,\delta$ are not IID. Absolute values are autocorrelated. This feature is also true for S&P 500 monthly returns themselves, see our recent article [17]. Also, these returns are not Gaussian. We wish to improve the fit of regressions (6) and (7) to make residuals closer to IID Gaussian. For S&P 500 returns, we did this in [17] as follows: We divided these returns by monthly average VIX: The S&P 500 Volatility Index $V$ computed daily by the Chicago Board of Options Exchange. Our main idea of this article [17] is (in our notation, see [17, Subsection 2.2]):
@@
This model (12) has good fit, as shown in [17, Section 3]. Innovations $W$ are IID but not Gaussian, with some finite exponential moments. The point estimate $\hat{\beta}=0.88$, and we reject the unit root hypothesis $\beta=1$.
@@ 2.5. Capital distribution curve
The plot $(\ln n,\ln\mu_{(n)}(t)),\,n=1,\ldots,N$ is called capital distribution curve. Empirically, it is a straight line for most $n$. If the model is stable, the capital distribution curve is also stable. Previous research was able to reproduce model stability and capital distribution curve for various continuous-time models: [6, 15]. These models were designed to capture the observations that, on average, small stocks have higher volatility and growth rates than large stocks. In [6], competing Brownian particle models were analyzed, where drift and diffusion coefficients depend on the rank of the stock. In [15], the drift and diffusion depend on the market weight of a stock. These classes of models do not use CAPM or VIX.
In the previous article by the second coauthor [11], devoted to the CAPM with size but without VIX, we also reproduced model stability for its continuous-time version [11, Theorem 2]. A natural question is to reproduce these results for our model here.
@@ 2.6. Our contributions
In [11], we model (7) and (6) but with $P_{0}$ or $S_{0}$ i.i.d. Gaussian, in continuous time. The article [11] does not include VIX $V$. However, this article is concerned with the idea from [17] that division by VIX normalized returns and premia. We replace $P$ and $P_{0}$ with $P/V$ and $P_{0}/V$ in (6). Also, we replace $R$ with $R/V$ and $R_{0}$ with $R_{0}/V$ in (7). We model $(R_{0}/V,P_{0}/V)$ as IID (but maybe not Gaussian), following (10) and (11).
@@ 3.3. Equity premia results
and $B$ is between $0.1$ and $0.2$ (except Decile 9, where we fail to reject $B=0$). Similarly to (13), this model is an improvement over [11] where residuals were not IID.
@@ 4.2. Stability results
The following is the main result of this article.
@@
Step 4. Finally, let us show ergodicity for the completed model. The transition function $\mathcal{Q}$ of this five-dimensional Markov chain is strictly positive: For any set $E\subseteq\mathbb{R}^{5}$ of positive Lebesgue measure, the transition probability $\mathcal{Q}(\mathbf{x},E)>0$ for every $\mathbf{x}\in\mathbb{R}^{5}$. Indeed, this transition probability is a push-forward of the distribution of the innovations in $\mathbb{R}^{d}$ under a certain smooth bijection $\mathbb{R}^{5}\to\mathbb{R}^{5}$. And for every $\mathbf{x}\in\mathbb{R}^{5}$, this probability measure $\mathcal{Q}(\mathbf{x},\cdot)$ and the Lebesgue measure on $\mathbb{R}^{5}$ are mutually absolutely continuous. For the rest of this proof, we refer the reader to the classic book [13] for terminology. It is straightforward to show that this Markov chain is irreducible and aperiodic. We have already shown it has a stationary distribution. Therefore, this Markov chain is positive Harris recurrent, and by [13, Theorem 13.0.1], this Markov chain is ergodic.
@@ 4.3. First-order approximation
The statement $\mathbb{E}[V^{u}(t)]<\infty$ for all $u\in(0,2]$ follows from the proof of [17, Lemma 1]. There, we can find that for MGF $M_{W}(u):=\mathbb{E}[e^{uW(t)}]$ we have:
@@ 5. Stochastic Portfolio Theory
Market weights and portfolios based on them are the main topic of Stochastic Portfolio Theory in both discrete time, see [5, 16, 20] and continuous time, see [10].
@@
Define by $\delta_{k}$ and $\varepsilon_{k}$ the sequences of innovations for equity premium, as in (18), and price returns, as in (19), for the $k$th portfolio, $k=1,\ldots,N$. Then the $32N+$ 3-dimensional vectors
@@
$C_{k}(t)=\ln\frac{S_{k}(t)}{S_{0}(t)},\quad k=1,\ldots,N.$
@@
This plot is called the capital distribution curve. With real-world markets, this curve is linear on most span, and concave overall. In our previous article [11], we captured the observation that well-diversified portfolios of small stocks have higher risk but higher return than that of large stocks. We reproduced this shape of capital distribution curve in [11] using simulation.
@@
The linearity of the capital distribution curve was rigorously proved in [6] for competing Brownian particles and in [15] for volatility-stabilized models. These two types of continuous-time models both capture this property (but without using CAPM) of small stocks.

### v3

@@ Abstract.
The Capital Asset Pricing Model (CAPM) relates a well-diversified stock portfolio to a benchmark portfolio. We insert size effect in CAPM, capturing the observation that small stocks have higher risk and return than large stocks, on average. Dividing stock index returns by the Volatility Index makes them independent and normal. In this article, we combine these ideas to create a new discrete-time model, which includes volatility, relative size, and CAPM. We fit this model using real-world data, prove the long-term stability, and connect this research to Stochastic Portfolio Theory. We fill important gaps in our previous article on CAPM with the size factor.
@@ 1. Introduction
Idea 2: Size Effect. Small stocks, on average, have higher volatility but also higher returns than large stocks. This feature implies long-term stability: Well-diversified stock portfolios stay together and not split into several clouds. We combined these two ideas in the previous article by the second author [13]. This article is a sequel of that work. Such long-term stability is of interest in Stochastic Portfolio Theory, which constructs portfolios as functions of market weights, utilizing the observation that small stocks have higher risk and return.
@@
In this article, we continue research of [13] by inserting Idea 3 in this setting, thus combining all three ideas. We create a new discrete-time model, and fit it using real market data from Kenneth French’s data library and Federal Reserve Economic Data web site. We then state and prove long-term stability of this market model, and interpret this for Stochastic Portfolio Theory. Unlike [13], we consider only discrete time models in this article. We fill a couple of important gaps left in the previous research [13].
In Section 2, we provide a comprehensive motivation of proposed models and historical review. Section 3 is devoted to data description and statistical fit of these models. In Section 4, we state and prove long-term stability (ergodicity) result for this model: Theorem 1, and analyze sufficient conditions for Theorem 1 to hold. We reduce the capital distribution curve to the order statistics (sorted values) of a standard normal sample. The last part of this section contains a short discussion of stability conditions for the case where volatility is constant. We have already discussed this in our previous article [13] but only for continuous time. Thus we fill a gap in our research.
In Section 5, we interpret our results in terms of Stochastic Portfolio Theory. We simulate capital distribution curve (ranked market weights vs ranks). We prove stability of this curve in Theorem 2. Next, Theorem 3 contains results about its shape: We reduce it to normal order statistics. This includes the case of constant volatility, which was studied in [13]. There, we did not have rigorous results; here, we fill this gap. Finally, the Appendix contains a discussion of the capital distribution curve based on ranked standard normal sample, continuing research of Section 5.
@@ 2.2. Size and value
Including a few factors (for example size and value) would enrich (3). In particular, size factor is related to $\beta$ as follows: Well-diversified portfolios of small stocks have equity premia $P$ closely correlated with that $P_{0}$ of large stock benchmark S&P 500. This simple linear regression of $P$ vs $P_{0}$ has very large $R^{2}$ but $\beta$ which is slightly larger as $1$. This can be done, for example, with exchange-traded funds tracking small, mid, and large (=benchmark) stock indices, see [13, Appendix]. The $\beta$ for mid-cap index is 1.15, and for small-cap index is 1.27. A natural idea then is to model $\beta$ as a function of a portfolio size relative to benchmark. We can try the same for $\alpha$, although in [13, Appendix] we have $\alpha$ is not significantly different from zero for both small-cap and mid-cap indices. See more on the size effect in [24], and the discussion in [1, Chapter 8].
@@ 2.3. CAPM-based model with size as factor
In fact, in [13] we have more general conditions: $\alpha$ and $\beta$ are general functions of $C$. Here, we focus only on linear functions, see [13, Example 1]. Analysis for small-cap and mid-cap funds above shows that $a\approx 0$ but $b<0$. Rewrite (3) as follows by plugging there (5):
@@ 2.4. Stochastic volatility
However, there is a drawback in our modeling from [13]. We fit regressions (6) and (7) using real-world monthly data from 1926 taken from Kenneth French’s Dartmouth College Financial Data Library. Residuals $\varepsilon,\delta$ are not IID. Absolute values are autocorrelated. This feature is also true for S&P 500 monthly returns themselves, see our recent article [22]. Also, these returns are not Gaussian. We wish to improve the fit of regressions (6) and (7) to make residuals closer to IID Gaussian. For S&P 500 returns, we did this in [22] as follows: We divided these returns by monthly average VIX: The S&P 500 Volatility Index $V$ computed daily by the Chicago Board of Options Exchange. Our main idea of this article [22] is (in our notation, see [22, Subsection 2.2]):
@@
We did not do this in [22], but we complete this work in this article. For the original equity premia of benchmark $P_{0}$, we plot in Figure 1 the quantile-quantile plot and the autocorrelation function for $P_{0}$ and $|P_{0}|$. For the normalized equity premia $P_{0}/V$, we plot these in Figure 2. We see that division by VIX is needed to model equity premia as in (11). Data for $P_{0}$ is taken from Kenneth French’s data library: Top 10% decile, January 1986–October 2024. For short-term rate $r$, we use the 3-month Treasury rate from Federal Reserve Economic Data: start of month data. Data and Python code is available on Github repository: asarantsev/size-capm-vix
(a) QQ
(b) ACF
(c) ACF
Figure 1. The quantile-quantile (QQ) plot for equity premia $P_{0}$, and the empirical autocorrelation function (ACF) for $P_{0}$ and for $|P_{0}|$
(a) QQ
(b) ACF
(c) ACF
Figure 2. The quantile-quantile (QQ) plot for equity premia $P_{0}/V$, and the empirical autocorrelation function (ACF) for $P_{0}/V$ and for $|P_{0}/V|$
@@
Slightly abusing the notation, but following [22], in the rest of the article we use $\alpha$ and $\beta$ as the intercept and the slope of this autoregression (12), instead of excess return and market exposure from the CAPM. This model (12) has good fit, as shown in [22, Section 3]. Innovations $W$ are IID but not Gaussian, with some finite exponential moments. The point estimate $\hat{\beta}=0.88$, and we reject the unit root hypothesis $\beta=1$.
@@ 2.5. Capital distribution curve
The plot of $N$ points
$(\ln n,\ln\mu_{(n)}(t)),\,n=1,\ldots,N$
is called the capital distribution curve. For real-world markets, this curve is concave and straight at left upper end, see [11, 12]. See also our own plots in Figure 3. Python code and data are available in GitHub repository asarantsev/size-capm-vix The data is from Kenneth French’s Data Library.
Figure 3. Split the US stock market into 20 equal parts based on market size: top 5%, next 5%, etc up to bottom 5%. Take the 19 breakpoints between these parts and use these as a substitution for market size $S_{k}$. We can compute market weights $\mu_{k}$ from these quantities $S_{k}$, for $k=1,\ldots,19$. We plot only 5 curves here for clarity but other months and years give very similar curves.
If the model is stable, the capital distribution curve is also stable in the long run: see Theorem 2. Previous research analyzed long-term stability and capital distribution curve for various continuous-time models: [7, 20]. These models were designed to capture the observations that, on average, small stocks have higher volatility and growth rates than large stocks. In [7], competing Brownian particle models were analyzed, where drift and diffusion coefficients depend on the rank of the stock, and linearity of the capital distribution curve was reproduced. In [20], the drift and diffusion depend on the market weight of a stock, but the capital distribution curve is not linear. These models do not use CAPM or VIX.
In the previous article by the second coauthor [13], devoted to the CAPM with size but without VIX (in other words, with constant volatility), we also reproduced model stability for its continuous-time version [13, Theorem 2]. Using simulation, we reproduced the linearity of the capital distribution curve. But we did not state and prove rigorous results on this. A natural question is to reproduce these results for our model here. We accomplish both tasks in this article: (a) we simulate the capital distribution curve in Subsection 5.2, which reproduces its linearity; (b) we prove results on convergence to a Poisson point process. This fills a gap in [13].
@@ 2.6. Our contributions
In [13], we model (7) and (6) but with $P_{0}$ or $S_{0}$ i.i.d. Gaussian, in continuous time. The article [13] does not include VIX $V$. However, this article is concerned with the idea from [22] that division by VIX normalized returns and premia. We replace $P$ and $P_{0}$ with $P/V$ and $P_{0}/V$ in (6). Also, we replace $R$ with $R/V$ and $R_{0}$ with $R_{0}/V$ in (7). We model $(R_{0}/V,P_{0}/V)$ as IID (but maybe not Gaussian), following (10) and (11). As mentioned in the Introduction, in Section 3 we perform statistical data analysis.
We fill two lacunas left in our research [13]. The first lacuna is stability results for discrete time. In Section 4 of this article, we state and prove this for the case of stochastic VIX. Our results work for constant volatility as well, which is the setting of [13], see (9). Next, we check the stability condition numerically. The second lacuna is rigorous results for the capital distribution curve, which are lacking in [13]. There, we present only the simulation for the constant volatility. In Section 5 of this article, we state and prove a convergence result for the capital distribution curve: Theorem 2. Next, in Theorem 3, we show a remarkable result: the curve is
$(\ln n,X_{(n)}),\quad n=1,\ldots,N,$
where $X_{(1)}\geq\ldots\geq X_{(N)}$ are order statistics of a conditional normal sample:
$X_{1},\ldots,X_{N}\mid\mu,\sigma\sim\mathcal{N}(\mu,\sigma^{2})$
for random $\mu$ and $\sigma$. We perform the simulation in Section 5 for stochastic VIX. In Appendix, we further analyze this curve for $\mu=$ and $\sigma=1$, and show its upper left and lower right ends replicate real-world behavior.
@@ 3.3. Equity premia results
and $B$ is between $-0.1$ and $-0.2$ (except Decile 9, where we fail to reject $B=0$). Similarly to (13), this model is an improvement over [13] where residuals were not IID.
@@ 4.2. Stability results
Below, we show that, if $\beta\in(0,1)$, these stationary versions $(V(t),R_{0}(t))$ exist, and Assumption 23 is well-defined. The following is the main result of this article.
@@
Step 4. Finally, let us show ergodicity for the completed model. The transition function $\mathcal{Q}$ of this five-dimensional Markov chain is strictly positive: For any set $E\subseteq\mathbb{R}^{5}$ of positive Lebesgue measure, the transition probability $\mathcal{Q}(\mathbf{x},E)>0$ for every $\mathbf{x}\in\mathbb{R}^{5}$. Indeed, this transition probability is a push-forward of the distribution of the innovations in $\mathbb{R}^{d}$ under a certain smooth bijection $\mathbb{R}^{5}\to\mathbb{R}^{5}$. And for every $\mathbf{x}\in\mathbb{R}^{5}$, this probability measure $\mathcal{Q}(\mathbf{x},\cdot)$ and the Lebesgue measure on $\mathbb{R}^{5}$ are mutually absolutely continuous. For the rest of this proof, we refer the reader to the classic book [17] for terminology. It is straightforward to show that this Markov chain is irreducible and aperiodic. We have already shown it has a stationary distribution. Therefore, this Markov chain is positive Harris recurrent, and by [17, Theorem 13.0.1], this Markov chain is ergodic.
@@ 4.3. First-order approximation
The statement $\mathbb{E}[V^{u}(t)]<\infty$ for all $u\in(0,2]$ follows from the proof of [22, Lemma 1]. There, we can find that for MGF $M_{W}(u):=\mathbb{E}[e^{uW(t)}]$ we have:
@@
4.4. The case of constant volatility
We stated and proved in [13] results on stationarity and ergodicity of a continuous-time analogue of our model, but with constant volatility. We did not state and prove stability result for discrete time. Here, we fill this gap. In the notation of this article, we can assume $V(t)=1$ is constant. This corresponds to $W(t)\equiv 0$.
Assume $Z(t)\sim\mathcal{N}(0,\sigma^{2})$ IID, and $V(t)=1$. This violates Assumption 1 (positive density). But this does not affect the stationarity proof. Let us check Assumption (9). We have: $R_{0}(t)=g+Z(t)$. Then we can rewrite the left-hand side of Assumption 23:
(33) $\begin{split}\mathbb{E}&\left[\ln|\xi|\right]<0,\quad\xi:=1+a+bR_{0}(t)\sim\mathcal{N}(\mu,\rho^{2}),\\ \mu&:=1+a+bg,\quad\rho:=|b|\sigma.\end{split}$
is shown in Figure 4. Python code is on GitHub repository asarantsev/size-capm-vix in file gaussian-simulation.py
@@ 5.1. Capital distribution curve
Market weights and portfolios based on them are the main topic of Stochastic Portfolio Theory in both discrete time, see [6, 21, 28] and continuous time, see [11, 12].
Define by $\delta_{k}$ and $\varepsilon_{k}$ the sequences of innovations for equity premium, as in (18), and price returns, as in (19), for the $k$th portfolio, $k=1,\ldots,N$:
(36) $\begin{split}\frac{P_{k}(t)}{V(t)}&=M+AC_{k}(t)+(1+BC_{k}(t))\frac{P_{0}(t)}{V(t)}+\delta_{k}(t);\\ \frac{R_{k}(t)}{V(t)}&=m+aC_{k}(t)+(1+bC_{k}(t))\frac{R_{0}(t)}{V(t)}+\varepsilon_{k}(t);\\ R_{k}(t)&=\ln\frac{S_{k}(t)}{S_{k}(t-1)},\quad C_{k}(t)=\ln\frac{S_{k}(t)}{S_{0}(t)}.\end{split}$
Together with (12), (16), (17), this is a time series Markov model for $32N+$ 3 series.
@@
The $32N+$ 3-dimensional vectors
@@
(37) $C_{k}(t)=\ln\frac{S_{k}(t)}{S_{0}(t)},\quad k=1,\ldots,N.$
@@
This plot is called the capital distribution curve. With real-world markets, this curve is linear on most span, and concave overall. Moreover, it shows remarkable long-term stability. See the famous picture in [11, Chapter 4] for eight capital dsitribution curves at end of years 1929, 1939, …, 1999; see the same picture as [12, Figure 13.4].
@@
In our previous article [13], we captured the observation that well-diversified portfolios of small stocks have higher risk but higher return than that of large stocks. We reproduced this shape of capital distribution curve in [13] using simulation.
From (37) and (34), we get a simple expression of $\mu_{k}(t)$ from $C_{k}(t)$:
$\ln\mu_{k}(t)=C_{k}(t)+\ln S_{0}(t)-\ln(S_{0}(t)+\ldots+S_{N}(t)),\,k=1,\ldots,N.$
Therefore, ranking market weights $\mu_{k}(t),k=1,\ldots,N$ from top to bottom at any fixed time $t$ is equivalent to ranking relative size terms $C_{k}(t),k=1,\ldots,N$ from top to bottom: $C_{(1)}(t)\geq\ldots\geq C_{(N)}(t)$. Thus instead of plotting the (slightly modified) capital distribution curve $(\ln k,\ln\mu_{(k)}(t)),k=1,\ldots,N$, we can plot $(\ln k,C_{(k)}(t)),k=1,\ldots,N$.
(a) Simulation 1
(b) Simulation 2
Figure 5. Two simulations of ranked relative size terms $(\ln k,C_{(k)}),k=1,\ldots,N$. We pick $(c,\rho)$ to be $(0.1,0)$, $(0.2,0)$, $(0.1,-0.5)$, $(0.2,-0.5)$
@@
The linearity of the capital distribution curve was rigorously proved in [7] for competing Brownian particles and disproved in [20] for volatility-stabilized models. These two types of continuous-time models both capture the property that small stocks have higher risk but higher return than large stocks. But these models do not use CAPM.
5.2. Simulation study
Instead of ranking all market weights, we can rank only weights of all portfolios but the benchmark. This technique will miss one portfolio from the capital distribution curve, but will not influence the overall behavior of the said curve. Consider the combined model:
  • Volatility $V$ governed by (12), with innovations $W$ having the variance-gamma distribution as in [22, subsection 3.3]
  • Benchmark price returns $R_{0}$ governed as (16) with Gaussian $Z$ having mean and standard deviation as in [22, Table 2, Large Price]
  • $N$ portfolios with relative size $C_{k}$ as in (37), $k=1,\ldots,N$, each governed by (22) for $a=m=0$ and negative $b$ to be chosen, with innovations $\varepsilon_{k}(t)\sim\mathcal{N}(0,1)$ independent for each $k$
For simplicity, we write all equations with number coefficients, replacing $-b$ by $c$:
$\ln V(t)$ $=0.346+0.882\ln V(t-1)+W(t),$
$W(t)$ $=0.0621+0.0621\Gamma(t)+0.1392\sqrt{\Gamma(t)}Y(t),$
$Y(t)$ $\sim\mathcal{N}(0,1),\quad\Gamma(t)\ \mbox{is Gamma with shape}\ 1/0.6573,\ \mathbb{E}[\Gamma(t)]=1,$
$R_{0}(t)$ $=V(t)Z(t),\quad Z(t)\sim\mathcal{N}(0.062,0.202^{2}),\quad\mathrm{Corr}(Y(t),Z(t))=\rho,$
$C_{k}(t+1)$ $=C_{k}(t)(1-cR_{0}(t))+V(t)\varepsilon_{k}(t),\quad\varepsilon_{k}(t)\sim\mathcal{N}(0,1),\quad k=1,\ldots,N,$
$(Y(t),Z(t)),\Gamma(t),\varepsilon_{1}(t),\ldots,\varepsilon_{N}(t)\ \mbox{independent for fixed}\ t$
$\bigl{(}Y(t),Z(t),\Gamma(t),\varepsilon_{1}(t),\ldots,\varepsilon_{N}(t)\bigr{)},\ t=1,2,\ldots\quad\mbox{IID}.$
We pick $N=100$ and simulate these equations for $T=400$ time steps, starting from
$\ln V(0)=3,\quad C_{k}(0)=0,\,k=1,\ldots,N.$
We do this for $t=T$, to give enough time to converge to the stationary distribution. We see various slopes for different cases: $\rho=0$ or $\rho=-50\%$ (this latter correlation is found in [22, Table 2, Large Price]); and $c=0.1$ or $c=0.2$ (lower and upper bound for $-b$ found in Subsection 3.2). We perform two simulations for each case. Each simulation has the same values of $V$ and $R_{0}$. We show all four curves for each simulation in Figure 5. See the Python code in the file capital-distribution-simulation.png from the GitHub repository asarantsev/size-capm-vix
5.3. Rigorous results: reduction to normal order statistics
Fix a constant $k=1,2,\ldots$. Assume the market weights $\mu_{n}$, or, equivalently, relative size $C_{n}$ are in the stationary distribution, which (by Theorem 2) is limiting distribution. To stress this, we write $t=\infty$ for time argument. Thus we have the ranked (sorted, ordered from top to bottom) values of the relative size process in the stationary distribution:
$C_{(1)}(\infty)>\ldots>C_{(N)}(\infty).$
Here, $N$ is the overall number of portfolios (excluding the benchmark). We are interested in the joint distribution of these sorted relative size values.
Assume $\mathcal{Z}_{1},\ldots,\mathcal{Z}_{N}\sim\mathcal{N}(0,1)$ IID is the standard normal sample.
Theorem 3.
There exists random variables $\mathcal{M}$ and $\mathcal{S}>0$ independent of the point process $\mathcal{N}$ which are functions of two time series: $V$ and $Z$, such that in law,
$\left[\frac{C_{1}(\infty)-\mathcal{M}}{\mathcal{S}},\ldots,\frac{C_{N}(\infty)-\mathcal{M}}{\mathcal{S}}\right]=(\mathcal{Z}_{1},\ldots,\mathcal{Z}_{N}).$
Figure 6. Three simulations of the standard normal market curve $(\ln k,\mathcal{Z}_{(k)})$ for $k=1,\ldots,N$ if $\mathcal{Z}_{1},\ldots,\mathcal{Z}_{N}\sim\mathcal{N}(0,1)$ IID sample, for $N=100$.
Proof.
Fix a $k=1,\ldots,N$. Apply [5, (0.6)] to express the stationary distribution of the relative size process $C_{k}$ from (22): letting
(38) $\begin{split}A_{n}&:=1+aV(n)+bR_{0}(n)=1+aV(n)+bV(n)(g+Z(n)),\\ B_{n}&:=V(n)(m+\varepsilon_{k}(n)),\end{split}$
pick any $n=1,2,\ldots$ and get:
(39) $C_{k}(\infty):=B_{n}+A_{n}B_{n-1}+A_{n}A_{n-1}B_{n-2}+\ldots$
We assume $V(n)$ and $Z(n)$ are defined for all $n\in\mathbb{Z}$, not just $n\geq 0$. Assume $\sigma_{\varepsilon}$ is the standard deviation of innovations $\varepsilon_{k}$. From (38) and (39), given time series $V$ and $Z$, the stationary distribution of $C_{k}(\infty)$ is Gaussian with mean and variance
$\mathcal{M}$ $:=m\left[V(n)+A_{n}V(n-1)+A_{n}A_{n-1}V(n-2)+\ldots\right],$
$\mathcal{S}^{2}$ $:=\sigma^{2}_{\varepsilon}\left[V^{2}(n)+A_{n}^{2}V^{2}(n-1)+A_{n}^{2}A_{n-1}^{2}V^{2}(n-2)+\ldots\right].$
Given time series $V$ and $Z$, the random variables $C_{1}(\infty),\ldots,C_{N}(\infty)$ are independent. Thus their standardized versions are (conditionally on $V$ and $Z$) IID standard Gaussian:
(40) $\mathcal{Z}_{n}:=\frac{C_{n}(\infty)-\mathcal{M}}{\mathcal{S}},\quad n=1,\ldots,N.$
This motivates the following definition.
Definition 4.
We define the standard normal market curve
(41) $(\ln k,\mathcal{Z}_{(k)}),\quad k=1,\ldots,N,$
for standard normal sample $\mathcal{Z}_{1},\ldots,\mathcal{Z}_{N}$ and its ranked version $\mathcal{Z}_{(1)}\geq\ldots\geq\mathcal{Z}_{(N)}$.
We study this curve in Appendix. The significance of Theorem 3 is as follows. Assume we rank $\mathcal{Z}_{(1)}>\ldots>\mathcal{Z}_{(N)}$. An immiediate consequence of Theorem 3 is that the (slightly modified) capital distribution curve $(\ln k,C_{(k)})$ for $k=1,\ldots,N$ has almost the same shape as the standard normal market curve. The only difference is a shift and change in slope, both random but independent of the standard normal market slope itself. Multiplication by $\mathcal{S}$ preserves ordering, since $\mathcal{S}>0$. We plotted three simulations of such curve in Figure 6. The Python code for this simulation in Figure 6 is available on GitHub repository asarantsev/size-capm-vix, file standard-normal-curve.py
It reproduces the real-world shape of the curve, but up to constant but random change in shift and slope, governed by $\mathcal{M}$ and $\mathcal{S}$. This is why the difference in simulations in Figure 5 is greater than in Figure 6. Indeed, much of the difference between two windows in Figure 5 (for example, between the two curves with $b=-0.1$ and zero correlation) is accounted for by the difference in $\mathcal{M}$ and $\mathcal{S}$ from simulation to simulation.
Changing the parameters of this model also affects the shape of the capital distribution curve, through the slope $\mathcal{S}$. This accounts for the differences within each window from Figure 5.